# On The Topology of Polygonal Meshes

### v2

@@ Abstract
This paper is an introductory and informal exposition on the topology of polygonal meshes. We begin with a broad overview of topological notions and discuss how homeomorphisms, homotopy, and homology can be used to characterise topology. We move on to define polygonal meshes and make a distinction between intrinsic topology and extrinsic topology which depends on the embedding space. A distinction is also made between quantitative topological properties and qualitative properties. Next, we outline proofs of the Euler and the Euler-Poincaré formulas. The Betti numbers are then defined in terms of the Euler-Poincaré formula and other mesh statistics rather than as cardinalities of the homology groups which allows us to avoid abstract algebra. Finally, we discuss how it is possible to cut a polygonal mesh such that it becomes a topological disc.
@@ 1 Introduction
While there are many applications of topology, it can be a challenging topic unless one is in possession of a good mathematical background; in particular, a firm grasp of abstract algebra is needed for algebraic topology. In this document, we target applications in visual computing and sidestep some of the mathematical requirements by considering only the topology of 3D surfaces provided as polygonal meshes. Moreover, we derive all quantitative topological properties from the Euler-Poincaré formula. By limiting the scope in this way, we are able to arrive at formulas for key topological properties such as genus and Betti numbers without group theory [alexandroff2012introduction]. Since Betti numbers are almost always defined in algebraic terms, this document takes an unusual route which I hope is illuminating for readers who wish to get a grasp on topology without necessarily understanding homology.
For readers who do want to understand homology, the book by Edelsbrunner and Harer [edelsbrunner2010computational] is a good introduction that focuses on the computational aspects, and Giblin’s book [giblin2013graphs] is recommended as a more mathematical text. Readers interested in a deeper understanding of the Euler-Poincaré formula should consider Imre Lakatos’ extraordinary book “Proofs and Refutations” [lakatos1963proofs] which uses the formula merely as an example, but since it is the only example it is covered very deeply.
@@ 1.1 Why is topology hard?
Topological properties often feel familiar and appeal to our intuition, but complications arise when we make the notions precise. One common source of confusion is related to the very notion of a hole. Consider a plastic bag that has been sealed. We could cut a hole in the bag with a pair of scissors in order to reach the inside of the bag. Now, consider a donut. Clearly, the donut already has a hole, but it is a different type of hole from the one in the plastic bag. Holes of the donut variety are shown in Figure 1.
Another pernicious source of confusion is the fact that the same object can be understood in different ways, and these different ways may lead to different conclusions regarding its topology. This will be quite relevant in the following when we disentangle intrinsic topology which is indifferent to how the object is embedded in space and extrinsic topology which does take the spatial arrangement into account.
To give another example of the importance of the framing, consider an egg-shaped object that is hollow inside. We might consider this object to consist of two surfaces: the outer surface and the surface that encloses the interior void. However, if we restrict our topological analysis to surfaces only, there is no way to distinguish an object with an interior cavity from two objects where one just happens to be inside the other. Since we will only be dealing with polygonal meshes in the following, it is worthwhile to be quite clear that polygonal meshes are surfaces, and there is no way of telling (from the representation itself) whether two disjoint polygonal meshes are part of the boundary of the same solid.
@@ 2.1 Boundaries
Faces are 2-dimensional patches bounded by edges. Likewise, edges are 1-dimensional curves (usually considered line segments) bounded by two vertices. The boundary relationship is formalised using the operator, $\partial$, which maps a set of $n$-dimensional mesh elements to their boundary which is a set of (n1)(n-1)-dimensional mesh elements. The boundary of a face is a set of edges,
@@ 2.5 Intrinsic and Extrinsic Topology
  • If we are interested in the intrinsic topology, the geometry is not important; a mesh is simply a schematic diagram of the structure of a surface. We can use this schematic to realise the mesh in 3D, but the intrinsic topology is independent of the realisation.
  • If we study the extrinsic topology, the realisation is the polygonal mesh. According to this view, a polygonal mesh is simply a set of points in 3D, namely the union of the points in contained in the faces. The topology of the mesh is that of the point set.
It is important to be aware of this distinction since the extrinsic and the intrinsic topology do not have to agree. Specifically, $\mathcal{M}$ may intersect itself in the embedding space which would often cause its extrinsic topology to differ from the intrinsic. For instance, a mesh may be intrinsically manifold but two vertices or two edges may touch as shown in Figure 5. In this case, the mesh is extrinsically non-manifold despite being manifold from the intrinsic point of view.
@@ 2.6 Qualitative Topology
The topological invariants watertightness and orientality are often important in the context of polygonal meshes. We can tell that they are topological properties since they are unaffected by a deformation of the mesh, but unlike the number of holes they are either true of false for a given mesh and thus qualitative rather than quantitative.
@@ 2.6.2 Orientability
Our procedure for setting the orientation of a face is extrinsic because it depends on choosing a direction in the embedding space. However, the cyclic ordering of the edges of a face is an intrinsic property, and whether a consistent ordering of all faces can be achieved is likewise intrinsic.

### v3

@@ Abstract
This paper is an introductory and informal exposition on the topology of polygonal meshes. We begin with a broad overview of topological notions and discuss how homeomorphisms, homotopy, and homology can be used to characterise topology. We move on to define polygonal meshes and make a distinction between intrinsic topology and extrinsic topology which depends on the space in which the mesh is immersed. A distinction is also made between quantitative topological properties and qualitative properties. Next, we outline proofs of the Euler and the Euler-Poincaré formulas. The Betti numbers are then defined in terms of the Euler-Poincaré formula and other mesh statistics rather than as cardinalities of the homology groups which allows us to avoid abstract algebra. Finally, we discuss how it is possible to cut a polygonal mesh such that it becomes a topological disc.
@@ 1 Introduction
While there are many applications of topology, it can be a challenging topic unless one has a relatively strong mathematical background. Specifically, it is important to understand group theory [alexandroff2012introduction], which belongs to the field of abstract algebra, since the important Betti numbers are defined as cardinalities of homology groups. In this document, we sidestep some of the mathematical requirements by considering the topology of 3D surfaces only in the form of polygonal meshes and derive all quantitative topological properties from the Euler-Poincaré formula. By limiting the scope in this way, we arrive at formulas for key topological properties such as genus and Betti numbers without referring to abstract algebra. For readers who do want to understand homology, the book by Edelsbrunner and Harer [edelsbrunner2010computational] is a good introduction that focuses on the computational aspects, and Giblin’s book [giblin2013graphs] is recommended as a more mathematical text. Readers interested in a deeper understanding of the Euler-Poincaré formula should consider Imre Lakatos’ extraordinary book “Proofs and Refutations” [lakatos1963proofs] which uses the formula merely as an example, but since it is the only example it is covered very thoroughly.
@@ 1.1 Why is topology hard?
[…]
Another pernicious source of confusion is the fact that the same object can be understood in different ways, and these different ways may lead to different conclusions regarding its topology. This will be quite relevant in the following when we disentangle intrinsic topology which is indifferent to how the object is immersed in space and extrinsic topology which does take the spatial arrangement into account. To give another example of the importance of the framing, consider an egg-shaped object that is hollow inside. We might consider this object to consist of two surfaces: the outer surface and the surface that encloses the interior void. However, if we restrict our topological analysis to surfaces only, there is no way to distinguish an object with an interior cavity from two objects where one just happens to be inside the other. Since we will only be dealing with polygonal meshes in the following, it is worthwhile to be quite clear that polygonal meshes are surfaces, and there is no way of telling (from the representation itself) whether two disjoint polygonal meshes are part of the boundary of the same solid.
@@ 2.1 Boundaries
From a higher dimensional mesh element (e.g. a face) we can obtain the lower dimensional elements that bound it (e.g. a set of edges) by exploiting the boundary relationship. The boundary relationship is formalised using the operator, $\partial$, which maps a set of $n$-dimensional mesh elements to their boundary which is a set of (n1)(n-1)-dimensional mesh elements. The boundary of a face is a set of edges,
@@ 2.5 Intrinsic and Extrinsic Topology
  • If we are interested in the intrinsic topology, the geometry is not important; a mesh is simply a schematic diagram of the structure of a surface. Since this diagram could be realised in 3D its topology is well defined, and this is what we call the intrinsic topology.
  • If we study the extrinsic topology, the mesh is immersed in 3D and this immersion is the polygonal mesh. According to this view, a polygonal mesh is simply a set of points in 3D, namely the image of the mesh under the immersion, and the extrinsic topology is that of the image. Put differently, the extrinsic topology is that which is induced by the immersion.
It is important to be aware of this distinction since the extrinsic and the intrinsic topology do not have to agree. Specifically, $\mathcal{M}$ may intersect itself in the immersion which would often cause its extrinsic topology to differ from the intrinsic. For instance, a mesh may be intrinsically manifold but two vertices or two edges may touch as shown in Figure 5. In this case, the mesh is extrinsically non-manifold despite being manifold from the intrinsic point of view.
@@ 2.6 Qualitative Topology
The topological invariants watertightness and orientability are often important in the context of polygonal meshes. We can tell that they are topological properties since they are unaffected by a deformation of the mesh, but unlike the number of holes they are either true of false for a given mesh and thus qualitative rather than quantitative.
@@ 2.6.2 Orientability
Our procedure for setting the orientation of a face is extrinsic because it depends on choosing a direction in the space into which it is immersed. However, the cyclic ordering of the edges of a face is an intrinsic property, and whether a consistent ordering of all faces can be achieved is likewise intrinsic.
@@ 8 Conclusions
Acknowledgements
Thanks to David Brander for helpful comments on nomenclature.